How can we Factorize x^2 - 3xy + y^2?


1
Expert's answer
2020-11-19T16:46:15-0500

"x^2 -3xy+y^2 = \\
=x^2-2xy+y^2-xy = \\
=(x-y)^2 - (\sqrt{xy})^2=\\
=(x-y-\sqrt{xy})(x-y+\sqrt{xy})\\ 
answer:\ (x-y-\sqrt{xy})(x-y+\sqrt{xy})"